OpenStudy (anonymous):

lim (x to infinity)(square root of ((x^2)+1)-x))

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty} \sqrt{x^2+1}-x\]

OpenStudy (anonymous):

gimmick is multiply top and bottom by conjugate

OpenStudy (anonymous):

get \[\frac{x^2+1-x^2}{\sqrt{x^2+1}+x}\] \[\frac{1}{\sqrt{x^2+1}+x}\] take limit get zero

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@satellite73: Hey!! I have posted a question, I would appreciate that you have a look at it :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

k, as long as it has nothing to do with laplace transforms from which i know from nothing

OpenStudy (anonymous):

Agreed satellite^^

OpenStudy (anonymous):

Lol, No! It's a programming problem. And by the way, Laplace transform is pretty easy :D

myininaya (myininaya):

anwar i dont see your question but i am not good at programming i can look though

OpenStudy (anonymous):

i wish i had learned to program. i should spend some time reading about laplace transforms too, big caesura in my education. looks ok, lots of formulas

myininaya (myininaya):

\[\lim_{x \rightarrow \infty }\frac{1}{\sqrt{x^2+1}+x}*\frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}}\] \[\lim_{x \rightarrow \infty} \frac{\frac{1}{\sqrt{x^2}}}{\sqrt{1+\frac{1}{x^2}}+\frac{x}{\sqrt{x^2}}}\] |x|=x when x>0 \[\lim_{x \rightarrow \infty}\frac{\frac{1}{x}}{\sqrt{1+\frac{1}{x^2}}+\frac{x}{x}}=\frac{0}{\sqrt{1+0}+1}=0\]